OpenStudy (anonymous):

Solve the equation:

OpenStudy (anonymous):

\[5 \sec^22\theta=\tan2\theta + 9 \] Give all solutions between 0-180 degrees. Yet another one I am stuck at, any help appreciated.

OpenStudy (ajprincess):

\(1+\tan^2\theta=\sec^2\theta\)

OpenStudy (ajprincess):

\(5(1+\tan^22\theta)=\tan2\theta+9\)

OpenStudy (anonymous):

Thank you, will have a go and see if I can get it done now. Was using sec=1/cos earlier.

OpenStudy (anonymous):

Got it, thanks again @ajprincess. Hopefully I can pass my finals next week :)

OpenStudy (ajprincess):

u r welcome. best of luck for ur finals.:)

OpenStudy (anonymous):

\[\cos^2(\theta) + \sin^2(\theta) = 1\] Divide by \[\cos^2(2\theta)\] we get \[1 + \tan^2(2\theta) = \sec^2(2\theta)\] With this result, plug it in your equation: \[5\left( 1+ \tan^2(2\theta) \right) = \tan(2\theta) + 9\] then \[5\tan^2(2\theta) - \tan(2\theta) - 4 = 0\] That's a quadratic equation on \[\tan(2\theta)\], just solve it: \[\tan(2\theta) = -4/5\] or \[\tan(2\theta) = 1\] for the first one: \[\theta = \frac{ \tan^{-1} (-4/5) }{ 2 } = -19.33º\] for the second one: \[\theta = \frac{ \tan^{-1} (1) }{ 2 } = (1/8)(4\pi n + \pi)\] Set n = 0 and n = 1: \[\theta = 22.5º\] and \[\theta = 112.5º\]